Unpolarized light of intensity 12 W/m^2 passes through a first polaroid filter and subsequently passes through a second filter with axes rotated by theta = 52 degrees relative to the first filter. What is the intensity of the light finally making it through both filters?

Question options:

12 W/m^2

6.0 W/m^2

4.5 W/m^2

2.3 W/m^2

9.3 W/m^2

Question

What is the strength of the electric field at the exact midpoint between charges +5.0 nC and -5.0 nC which are separated by 10. cm?

Question options:

1.8 x 10^4 N/C

3.6 x 10^4 N/C

none of these

9.0 x 10^3 N/c

7.2 x 10^4 N/C

Question

A skydiver is moving toward the Earth at 56 m/s (terminal velocity) while holding a loud tone generator producing sinusoidal sound at 1.2 kHz.   Her friend is on the ground observing everything.   What frequency does the stationary (and safe) observer register on the ground? [Take the speed of sound to be 343 m/s].

Question options:

1.4 kHz

1.6 kHz

1.8 kHz

none of these

1.2 kHz

Question

Charges q_1 = 2.00 micro C and q_2 = 200. micro C are nearby one another. Which one exerts a greater force on the other?

Question options:

q_1 and q_2 exert equal but opposite forces on one another

q_1 and q_2 exert exactly the same force on one another

q_2 exerts the greater force

q_1 exerts the greater force

the answer depends upon how far apart they are
Answer #1

0 Om One Iz I= 12 W/m2 I = 77 / 2 = 12 / 2 = 6 W/m2 I = I, cos 0 = 6 losse = 6x(0.615) - 2.27 W/m² = 2:3 E = 2 k2 2 = 2X9 x 1